# Electron-phonon coupling induced topological phase transition in an $\alpha$-$T_{3}$ Haldane-Holstein model

### v1

@@ III.3 Berry curvature and Chern number
To study the topological signatures in the $\alpha$-$T_{3}$ induced by the e-ph coupling, we numerically compute the topological ingredients, namely the (polaronic) Berry curvature and the Chern number. We also obtain the phase diagram containing the Chern number and e-ph coupling strength. In a usual $\alpha$-$T_{3}$ lattice, due to the TRS breaking NNN Haldane term, the system exhibits a nonzero Chern number. The onsite Samenoff mass term that breaks the valley degeneracy also plays a crucial role in band opening at high symmetry Dirac points.
Figure 10: Berry curvature corresponding to the VB is presented for lower $\alpha$-values in different regimes of $\lambda_{eph}$ as mentioned in Fig. 3.
Figure 11: Berry curvature corresponding to the VB is presented for $\alpha=0.7$ in different regimes of $\lambda_{eph}$ as mentioned in Fig. 5.
Figure 12: Berry curvature corresponding to the VB is presented for $\alpha=0.8$ in different regimes of $\lambda_{eph}$ as mentioned in Fig. 6.
@@
To show the variations of the Berry curvature of $\alpha=0.7$, we plot Fig. 11 that can be explained with the help of the bulk and edge spectra displayed in Fig. 4 and Fig. 8, respectively. Let us first look at Figs. 11 $(a)$ and $(b)$ which is for $\lambda_{eph}=0$ and $\lambda_{eph}<\lambda_{c_{1}}$, respectively. In this regime, as we have discussed, the spectral gap between the FB and VB vanishes even at $\lambda_{eph}=0$, and remains so till $\lambda_{eph}=\lambda_{c_{1}}$, manifesting the chiral edge states (see Fig. 8$(a)$) at the boundaries. However, in this regime, the Berry curvature shows singular behaviour as there is no bulk gap, and consequently, the Chern number is ill-defined. But as we tune $\lambda_{eph}$ further, a bulk gap opens up for the first time at around $\lambda_{eph}\sim\lambda_{c_{1}}$ (see Fig. 4)$(b)$), where we see that the concentrations at the K-points start behaving differently than that at $\bf{K^{\prime}}$-points and are shown in Fig. 11$(b)$. Beyond $\lambda_{c_{1}}$, this signature is much more noticeable (can be seen in Fig. 11$(c)$) and the edge states are prominent (see Fig. 8$(b)$) in $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime. In the vicinity of $\lambda_{eph}=\lambda_{c_{2}}$, the bulk gap closes showing high values for the Berry curvature (see Fig. 11$(d)$). Finally beyond $\lambda_{c_{2}}$ (Fig. 11$(e)$), the variation of the Berry curvature is reminiscent of Figs. 10 $(c)$, $(f)$ and $(i)$ enunciates the onset of a trivial insulating phase.
The Berry curvature plots for $\alpha=0.8$ are displayed in Fig. 12. Although the variations in the $\lambda_{eph}<\lambda_{c_{1}}$ regime (Fig. 12$(b)$) may look similar to those for $\alpha=0.7$ showing higher values of the Berry curvatures even for $\lambda_{eph}=0$ (Fig. 12$(a)$), but with the support of the findings of Fig. 6 and Fig. 9 described in Sec. III.1 and III.2 respectively, it is ensured that in the $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime, they may correspond to some topological phase (unlike the usual $SM$ phase for $\alpha=0.7$) with conducting edge modes (see Figs. 9$(a-b)$) associated with higher Chern numbers. The observations of Figs. 12$(c-f)$ are almost same as $\alpha=0.7$ case. However, a noticeable dissimilarity with the $\alpha=0.7$-variation in the Berry curvature can be observed for the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime, where the disparity between the concentrations of the Berry curvature at two valleys is much more significant compared to that for $\alpha=0.7$. This makes the variation of Fig. 12$(d)$ distinguishable from Fig. 12$(a)$ denoting different topological phases.
@@
Figure 14: The Chern number, $C$ corresponding to the VB as a function of e-ph coupling strength, $\lambda_{eph}$ for $(i)$ lower to intermediate $\alpha$ values ($0<\alpha\leq 0.6$) is shown, while in the inset $(a)$ a zoomed-in picture of the transition regions, and in the inset $(b)$, the variation of $\lambda_{c}$ as a function of $\alpha$ is shown. In $(ii)$, the variations of $C$ for larger $\alpha$ values ($0.6<\alpha\leq 0.9$) are shown. The inset $(a)$ represents a zoomed-in picture of the transition regions, while the insets $(b)$ and $(c)$, respectively, display the variations of $\lambda_{c_{2}}$ and $\lambda_{c_{1}}$ as a function of $\alpha$. The values of $\lambda_{c}$ are mentioned in Table 2 and 2.
@@
Now let us examine the dependency of the Chern number on $\alpha$ in the presence of the e-ph interaction. Fig. 14 displays the variations of the Chern number as a function of $\lambda_{eph}$. Here, we display the variations of $C$ separately in two diagrams for lower to intermediate values of $\alpha$ ($\alpha=0.1,0.2,...,0.6$) (see Fig. 14$(i)$) and larger values of $\alpha$ ($\alpha=0.7,0.8,0.9$) (see Fig. 14$(ii)$). Starting from the $\alpha=0.1$ case to the intermediate values, such as $\alpha=0.6$, we notice that $C=-1$ up to a critical $\lambda_{c}$, at which $C$ abruptly falls to $C=0$ showing a sharp discontinuity. Therefore, for lower to intermediate cases of $\alpha$ (Fig. 14$(i)$), the system initially behaves like a Chern insulator designated by a nonzero Chern number until the e-ph coupling reaches a certain critical value, $\lambda_{c}$ ($\lambda_{c}$’s are listed in Table 2) at which the system undergoes a topological transition accompanied by the closing of the bulk gap and emerging signatures of the edge states. While, beyond $\lambda_{c}$, the system ceases to host edge states which is a typical signature of a trivial insulator for which $C=0$. We display a zoomed-in picture of the transition points in the inset $(a)$ of Fig. 14$(i)$ and $(ii)$. In the inset $(b)$ of Fig. 14$(i)$, we depict the variation of the critical $\lambda_{c}$ with respect to $\alpha$, illustrating a nearly linear decrease with increasing $\alpha$.
@@
It is understood by now that such variations at higher $\alpha$ values are in contrast to those at lower values of $\alpha$. Let us first consider the variation corresponding to $\alpha=0.7$ which is represented via a solid blue line in Fig. 14$(ii)$. Unlike the lower $\alpha$ values, there exist two transition points, namely, $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ for higher $\alpha$ values (listed in Table 2). Below the former, the system inherits a conventional $SM$ phase, and above the latter, the system becomes a trivial insulator. Understandably, the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ region is our main interest for $\alpha=0.7$, where we find that the Chern number is fixed at $C=-1$ that underscores the emergence of a topologically nontrivial insulating phase, driven entirely by the e-ph coupling. As expected, beyond $\lambda_{c_{2}}$, $C$ becomes zero confirming the onset of a trivial phase. Thus, the e-ph coupling favours a transition from a semi-metal to a topological insulator, and to a trivial insulator for $\alpha=0.7$. However, the $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime becomes interesting for $\alpha>0.7$ as described earlier in the findings of Figs. 9 $(a)$ and $(b)$ that the conducting edge modes exist (specifically for $10.8\leq\alpha<$ 1) in that regime of $\lambda_{eph}$ indicating a topologically nontrivial insulating phase. In Fig. 14$(ii)$, we plot $C$ as a function of $\lambda_{eph}$ for $\alpha=0.8$ denoted by the solid orange line, where higher Chern number, namely $C=-2$ in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime is noted, confirming emergence of a distinct (other than $C=-1$) topological phase. Nevertheless, as we tune $\lambda_{eph}$ further, the scenario becomes exactly the same as $\alpha=0.7$, that is, $C$ changes from $C=-2$ to $C=-1$ at $\lambda_{eph}=\lambda_{c_{1}}$ signifying a different topological phase that persists in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime, which finally vanishes beyond $\lambda_{c_{2}}$. A similar observation is also shown for $\alpha=0.9$ (marked by the solid green line). Therefore, for the $10.7<\alpha<$ 1 regime, the system undergoes a transition from one topological phase ($C=-2$) to another ($C=-1$) and hence transits to a trivial ($C=0$) phase, purely mediated all the while by the e-ph coupling. The emergence of $|C|=2$ topological phase in an $\alpha-T_{3}$ is a familiar phenomenon obtained by others Wang and Liu (2021); Dey and Ghosh (2019); Mondal and Basu (2023) in the absence of e-ph interaction. For our case, the results completely match with those in Ref. Wang and Liu (2021) corresponding to $\lambda_{eph}\rightarrow 0$ and $M\rightarrow 0$ (shown in Fig.13). Moreover, due to the e-ph interaction, we obtain a $|C|=2$ topological phase for lower values of $\lambda_{eph}$, even at $\lambda_{eph}=0$ (in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime) for $10.8\leq\alpha<$ 1, that is for $\alpha$ values close to the dice lattice limit ($\alpha=1$). As earlier, in the inset $(b)$ and $(c)$ of Fig. 14$(ii)$, we show the variation of the critical $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ with respect to $\alpha$. We observe that $\lambda_{c_{1}}$ initially decreases and then increases with increasing $\alpha$, whereas an opposite trend is observed for $\lambda_{c_{2}}$, namely, it increases first and hence decreases with increasing $\alpha$. We should mention that the findings of Chern number plots are completely consistent with those of bulk and edge spectra for different regimes of $\lambda_{eph}$. As stated in Sec. III.1, although we have shown a few cases of $\alpha$, it is also important to note that this kind of transition can occur for any value of $\alpha$ ($10<\alpha<$ 1). Therefore, it seems robust that the polaron formation in $\alpha$-$T_{3}$ lattices induces a topological phase transition generated solely due to the presence of e-ph coupling.
@@
Figure 15: The topological phase diagram based on the Chern number ($C$) corresponding to the VB in $\lambda_{eph}-\alpha$ plane. The nonzero Chern number corresponding to the teal region is denoted as $C=-1$, while the yellow region represents the vanishing Chern number ($C=0$). The grey region denotes the $SM$ phase for $0.65\lesssim\alpha\lesssim 0.75$, while the deep purple region stands for a distinct topological phase with $C=-2$ in the regime $10.75\lesssim\alpha<$ 1. Other parameters remain the same as mentioned in Fig. 3.
@@ III.4 Hall conductivity
Figure 16: The Hall conductivity, $\sigma_{xy}$ as a function of Fermi energy, $E_{f}$ is presented for various values of $\alpha$, with $(a)$ $\alpha=0.1$, $(b)$ $\alpha=0.3$, $(c)$ $\alpha=0.7$, and $(d)$ $\alpha=0.8$ and $\lambda_{eph}$ values are shown in the inset. Other parameters are same as mentioned in Fig. 3.
@@
where $e^{2}/h=\sigma_{0}$ is the scale in which $\sigma_{xy}$ is measured, $E_{k_{x},k_{y}}^{\gamma}$ is the energy band with the band index $\gamma=-1,0$ and $+1$ corresponding to the valance, flat and conduction bands, respectively, $f$ denotes the Fermi-Dirac distribution function: $f(E)=[1+e^{(E-E_{F})/k_{B}T}]^{-1}$, $E_{F}$ and $T$ being the Fermi energy and the absolute temperature, respectively, $\Omega(k_{x},k_{y})$ being the Berry curvature. Fig. 16 display the variations of the polaronic Hall conductivities at $T=0$ as a function of $E_{f}$ for different values of e-ph interaction strength, $\lambda_{eph}$ for $\alpha=0.1$ (see Fig. 16$(a)$), $\alpha=0.3$ (see Fig. 16$(b)$), $\alpha=0.7$ (Fig. 16$(c)$), and $\alpha=0.8$ (Fig. 16$(d)$).
@@
As shown in Fig. 16$(a)$, the Hall conductivity ($\sigma_{xy}$) is plotted as a function of the Fermi energy ($E_{f}$) for $\lambda_{eph}=0.3,0.4$ and $0.5$ marked by solid blue, black and red colours, respectively. It is observed for $\alpha=0.1$, the Hall conductivities initially increase and show tiny plateaus (can also be seen in Fig. 16$(b)$ and Fig. 16$(c)$ for $\alpha=0.3$ and $\alpha=0.7$, respectively), which are quantized at a value $e^{2}/h$ for the $\lambda_{eph}<\lambda_{c}$ regime. In other words, these quantized plateaus occurring at $|C|e^{2}/h$ (here, $|C|=1$) presented in Fig. 16$(a)$ re-confirm that up to a critical e-ph coupling $\lambda_{c}\approx 0.49$, the system behaves like a topological insulator. Beyond the critical $\lambda_{c}$ (denoted by solid red), it becomes a trivial insulator with $\sigma_{xy}=0$ (that is, $C=0$). Similar observation is noted in Fig. 16$(c)$ for $\alpha=0.7$ where the plateaus at values $|C|e^{2}/h$ exist for $\lambda_{eph}$ values that are in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime.
@@
In a scenario where $\lambda_{eph}=0=M$, the Hall conductivity shows plateaus (but with small kinks on the plateaus due to the presence of the distorted FB) as long as the Fermi level lies in between the bulk gap Mondal and Basu (2023). However, in our case (with $\lambda_{eph}\neq 0$ and $M\neq 0$), the nature of the Hall conductivity deviates significantly as there exists a cumulative effect arising from the interplay of the three parameters, namely the Fermi energy ($E_{f}$), the mass term ($M$) and the e-ph coupling strength ($\lambda_{eph}$). The reason for the plateaus to become tinier can be explained with the help of the bulk spectra, which are mainly affected by $M$ and $\lambda_{eph}$ for different $\alpha$ values. It is understandable that a significant width of the plateau is dependent on how accurately we fix the Fermi level in the bulk gap. As discussed in Sec. III.1, the individual bulk bands shrink due to the Holstein factor (Eq. (8)), and the whole band structure shifts vertically down by the polaron shift energy ($\lambda_{eph}^{2}\hbar\omega_{0}$). Due to the band narrowing caused by the Holstein factor, the gap between the distorted (because of the Haldane term, $\lambda$) FB and VB decreases, makes it difficult for the Fermi level to lie ‘properly’ in between the bulk gap, making the plateaus less prominent, especially for higher values of $\lambda_{eph}$. Additionally, we find that the increase in $\sigma_{xy}$ as a function of $E_{f}$ can be explained as follows. As observed in Sec. III.1, $M$ breaks the valley degeneracy and the interplay between $M$ and $\lambda_{eph}$ renders contrasting behaviour of the bulk bands at two valleys, that is, well-gapped at one valley and almost gapless at the other for $\lambda_{eph}<\lambda_{c}$. As $E_{f}$ is increased, it is possible that at one valley, $E_{f}$ may lie well in the gap, while it may lie in the CB as well at the other valley, which will contribute to higher $\sigma_{xy}$. The unusual behaviour of the Hall conductivity due to the presence of a distorted FB has also been reported by Singh et al. Singh and Sharma (2023a). Certainly, all of the above discussions become unimportant for $\lambda_{eph}>\lambda_{c}$.

### v2

@@ III.3 Berry curvature and Chern number
To ascertain the topological signatures in the $\alpha$-$T_{3}$ induced by the e-ph coupling, we numerically compute the topological ingredients, namely the (polaronic) Berry curvature and the Chern number. We also obtain the phase diagram containing the Chern number and e-ph coupling strength. In a usual $\alpha$-$T_{3}$ lattice, due to the TRS breaking NNN Haldane term, the system exhibits a nonzero Chern number. The onsite Samenoff mass term that breaks the valley degeneracy also plays a crucial role in band opening at high symmetry Dirac points.
[…]
Figure 11: The Berry curvature corresponding to the VB is presented for $\alpha=0.7$ in different regimes of $\lambda_{eph}$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.2$), $(c)$ $\lambda_{eph}\sim\lambda_{c_{1}}$ ($\lambda_{eph}=0.26$), $(d)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(e)$ $\lambda_{eph}\sim\lambda_{c_{2}}$ ($\lambda_{eph}=0.41$), and $(f)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). Other parameters are mentioned in Fig. 5. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.
Figure 12: The Berry curvature corresponding to the VB is presented for $\alpha=0.8$ in different regimes of $\lambda_{eph}$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.15$), $(c)$ $\lambda_{eph}\sim\lambda_{c_{1}}$ ($\lambda_{eph}=0.18$), $(d)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(e)$ $\lambda_{eph}\sim\lambda_{c_{2}}$ ($\lambda_{eph}=0.42$), and $(f)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). Other parameters are mentioned in Fig. 6. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.
@@
To show the variations of the Berry curvature of $\alpha=0.7$, we plot Fig. 11 that can be explained with the help of the bulk and edge spectra displayed in Fig. 5 and Fig. 8, respectively. Let us first look at Figs. 11 $(a)$ and $(b)$ which is for $\lambda_{eph}=0$ and $\lambda_{eph}<\lambda_{c_{1}}$, respectively. In this regime, as we have discussed, the spectral gap between the FB and VB vanishes even at $\lambda_{eph}=0$, and remains so till $\lambda_{eph}=\lambda_{c_{1}}$, manifesting the chiral edge states (see Figs. 8$(a-b)$) at the boundaries. However, in this regime, the Berry curvature shows singular behaviour as there is no bulk gap, and consequently, the Chern number is ill-defined. But as we tune $\lambda_{eph}$ further, a bulk gap opens up for the first time at around $\lambda_{eph}\sim\lambda_{c_{1}}$ (see Fig. 5)$(c)$), where we see that the concentrations at the K-points start behaving differently than that at $\bf{K^{\prime}}$-points and are shown in Fig. 11$(c)$. Beyond $\lambda_{c_{1}}$, this signature is much more noticeable (can be seen in Fig. 11$(d)$) and the edge states are prominent (see Fig. 8$(d)$) in $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime. In the vicinity of $\lambda_{eph}=\lambda_{c_{2}}$, the bulk gap closes showing high values for the Berry curvature (see Fig. 11$(e)$). Finally beyond $\lambda_{c_{2}}$ (Fig. 11$(f)$), the variation of the Berry curvature is reminiscent of Figs. 10 $(c)$, $(f)$ and $(i)$ enunciates the onset of a trivial insulating phase.
The Berry curvature plots for $\alpha=0.8$ are displayed in Fig. 12. Although the variations in the $\lambda_{eph}<\lambda_{c_{1}}$ regime (Fig. 12$(b)$) may look similar to those for $\alpha=0.7$ showing higher values of the Berry curvatures even for $\lambda_{eph}=0$ (Fig. 12$(a)$), but with the support of the findings of Fig. 6 and Fig. 9 described in Sec. III.1 and III.2 respectively, it is ensured that in the $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime, they may correspond to some topological phase (unlike the usual $SM$ phase for $\alpha=0.7$) with conducting edge modes (see Figs. 9$(a-b)$) associated with higher Chern numbers. The observations of Figs. 12$(c-f)$ are almost same as $\alpha=0.7$ case. However, a noticeable dissimilarity with the $\alpha=0.7$-variation in the Berry curvature can be observed for the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime (Fig. 12$(d)$), where the disparity between the concentrations of the Berry curvature at two valleys is much more significant compared to that for $\alpha=0.7$. This makes the variation of Fig. 12$(d)$ distinguishable from Figs. 12$(a-b)$ denoting different topological phases.
@@
Figure 14: The Chern number, $C$ corresponding to the VB as a function of e-ph coupling strength, $\lambda_{eph}$ for $(i)$ lower to intermediate $\alpha$ values ($0<\alpha\leq 0.6$) is shown, while in the inset $(a)$ a zoomed in picture of the transition regions, and in the inset $(b)$, the variation of $\lambda_{c}$ as a function of $\alpha$ is shown. In $(ii)$, the variations of $C$ for larger $\alpha$ values ($0.6<\alpha\leq 0.9$) are shown. The inset $(a)$ represents a zoomed in picture of the transition regions, while the insets $(b)$ and $(c)$, respectively, display the variations of $\lambda_{c_{2}}$ and $\lambda_{c_{1}}$ as a function of $\alpha$. The values of $\lambda_{c}$ are mentioned in Table 2 and 2.
@@
Now let us examine the dependency of the Chern number on $\alpha$ in the presence of the e-ph interaction. Fig. 14 displays the variations of the Chern number as a function of $\lambda_{eph}$. Here, we display the variations of $C$ separately in two diagrams for lower to intermediate values of $\alpha$ ($\alpha=0.1,0.2,...,0.6$) (see Fig. 14$(i)$) and larger values of $\alpha$ ($\alpha=0.7,0.8,0.9$) (see Fig. 14$(ii)$). Starting from the $\alpha=0.1$ case to the intermediate values, such as $\alpha=0.6$, we notice that $C=-1$ up to a critical $\lambda_{c}$, at which $C$ abruptly falls to $C=0$ showing a sharp discontinuity. Therefore, for lower to intermediate cases of $\alpha$ (Fig. 14$(i)$), the system initially behaves like a Chern insulator designated by a nonzero Chern number until the e-ph coupling reaches a certain critical value, $\lambda_{c}$ ($\lambda_{c}$’s are listed in Table 2) at which the system undergoes a topological transition accompanied by the closing of the bulk gap and emerging signatures of the edge states. While, beyond $\lambda_{c}$, the system ceases to host edge states which is a typical signature of a trivial insulator for which $C=0$. We display a zoomed in picture of the transition points in the inset $(a)$ of Fig. 14$(i)$ and $(ii)$. In the inset $(b)$ of Fig. 14$(i)$, we depict the variation of the critical $\lambda_{c}$ with respect to $\alpha$, illustrating a nearly linear decrease with increasing $\alpha$.
@@
It is understood by now that such variations at higher $\alpha$ values are in contrast to those at lower values of $\alpha$. Let us first consider the variation corresponding to $\alpha=0.7$ which is represented via a solid blue line in Fig. 14$(ii)$. Unlike the lower $\alpha$ values, there exist two transition points, namely, $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ for higher $\alpha$ values (listed in Table 2). Below the former, the system inherits a conventional $SM$ phase (where $C$ is ill-defined), and above the latter, the system becomes a trivial insulator. Understandably, the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ region is our main interest for $\alpha=0.7$, where we find that the Chern number is fixed at $C=-1$ that underscores the emergence of a topologically nontrivial insulating phase, driven entirely by the e-ph coupling. As expected, beyond $\lambda_{c_{2}}$, $C$ becomes zero confirming the onset of a trivial phase. Thus, the e-ph coupling favours a transition from a semi-metal to a topological insulator, and to a trivial insulator for $\alpha=0.7$. However, the $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime becomes interesting for $\alpha>0.7$ as described earlier in the findings of Figs. 9 $(a)$ and $(b)$ that the conducting edge modes exist (specifically for $10.8\leq\alpha<$ 1) in that regime of $\lambda_{eph}$ indicating a topologically nontrivial insulating phase. In Fig. 14$(ii)$, we plot $C$ as a function of $\lambda_{eph}$ for $\alpha=0.8$ denoted by the solid orange line, where higher Chern number, namely $C=-2$ in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime is noted, confirming emergence of a distinct (other than $C=-1$) topological phase. Nevertheless, as we tune $\lambda_{eph}$ further, the scenario becomes exactly the same as $\alpha=0.7$, that is, $C$ changes from $C=-2$ to $C=-1$ at $\lambda_{eph}=\lambda_{c_{1}}$ signifying a different topological phase that persists in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime, which finally vanishes beyond $\lambda_{c_{2}}$. A similar observation is also shown for $\alpha=0.9$ (marked by the solid green line). Therefore, for the $10.7<\alpha<$ 1 regime, the system undergoes a transition from one topological phase ($C=-2$) to another ($C=-1$) and hence transits to a trivial ($C=0$) phase, purely mediated all the while by the e-ph coupling. The emergence of $|C|=2$ topological phase in an $\alpha$-$T_{3}$ is a familiar phenomenon obtained by others Wang and Liu (2021); Dey and Ghosh (2019); Mondal and Basu (2023) in the absence of e-ph interaction. For our case, the results completely match with those in Ref. Wang and Liu (2021) corresponding to $\lambda_{eph}\rightarrow 0$ and $M\rightarrow 0$ (shown in Fig.13). Moreover, due to the e-ph interaction, we obtain a $|C|=2$ topological phase for lower values of $\lambda_{eph}$, even at $\lambda_{eph}=0$ (in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime) for $10.8\leq\alpha<$ 1, that is for $\alpha$ values close to the dice lattice limit ($\alpha=1$). As earlier, in the inset $(b)$ and $(c)$ of Fig. 14$(ii)$, we show the variation of the critical $\lambda_{c_{2}}$ and $\lambda_{c_{1}}$, respectively as a function of $\alpha$. We observe that $\lambda_{c_{1}}$ initially decreases and then increases with increasing $\alpha$, whereas an opposite trend is observed for $\lambda_{c_{2}}$, namely, it increases first and hence decreases with increasing $\alpha$. We should mention that the findings of Chern number plots are completely consistent with those of bulk and edge spectra for different regimes of $\lambda_{eph}$. As stated in Sec. III.1, although we have shown a few cases of $\alpha$, it is also important to note that this kind of transition can occur for any value of $\alpha$ ($10<\alpha<$ 1). Therefore, it seems robust that the polaron formation in $\alpha$-$T_{3}$ lattices induces a topological phase transition generated solely due to the presence of e-ph coupling.
@@
Figure 15: The topological phase diagram based on the Chern number ($C$) corresponding to the VB in $\lambda_{eph}-\alpha$ plane. The nonzero $C$ corresponding to the teal region is denoted as $C=-1$, while the yellow region represents the vanishing Chern number ($C=0$). The grey region denotes the $SM$ phase for $0.65\lesssim\alpha\lesssim 0.75$, while the deep purple region stands for a distinct topological phase with $C=-2$ for $10.75\lesssim\alpha<$ 1. Other parameters remain the same as mentioned in Fig. 3.
@@ III.4 Hall conductivity
Figure 16: The Hall conductivity, $\sigma_{xy}$ as a function of Fermi energy, $E_{f}$ is presented for various values of $\alpha$: $(a)$ $\alpha=0.1$, $(b)$ $\alpha=0.3$, $(c)$ $\alpha=0.7$, and $(d)$ $\alpha=0.8$ for different $\lambda_{eph}$ values that are shown in the inset. Other parameters are the same as mentioned in Fig. 3.
@@
where $e^{2}/h=\sigma_{0}$ is the scale in which $\sigma_{xy}$ is measured, $E_{k_{x},k_{y}}^{\gamma}$ is the energy band with the band index $\gamma=-1,0$ and $+1$ corresponding to the VB, FB and the CB, respectively, $f$ denotes the Fermi-Dirac distribution function: $f(E)=[1+e^{(E-E_{F})/k_{B}T}]^{-1}$, $E_{F}$ and $T$ being the Fermi energy and the absolute temperature, respectively, $\Omega(k_{x},k_{y})$ being the Berry curvature. Fig. 16 display the variations of the polaronic Hall conductivities at $T=0$ as a function of $E_{f}$ for different values of e-ph interaction strength, $\lambda_{eph}$ for $\alpha=0.1$ (see Fig. 16$(a)$), $\alpha=0.3$ (Fig. 16$(b)$), $\alpha=0.7$ (Fig. 16$(c)$), and $\alpha=0.8$ (Fig. 16$(d)$).
@@
As shown in Fig. 16$(a)$, the Hall conductivity ($\sigma_{xy}$) is plotted as a function of the Fermi energy ($E_{f}$) for $\lambda_{eph}=0.3,0.4$ and $0.5$ marked by solid blue, black and red colours, respectively. It is observed for $\alpha=0.1$, the Hall conductivities initially increase and show tiny plateaus (can also be seen in Fig. 16$(b)$ and Fig. 16$(c)$ for $\alpha=0.3$ and $\alpha=0.7$, respectively), which are quantized at a value $e^{2}/h$ for the $\lambda_{eph}<\lambda_{c}$ regime. In other words, these quantized plateaus occurring at $|C|e^{2}/h$ (here, $|C|=1$) presented in Fig. 16$(a)$ re-confirm that up to a critical e-ph coupling $\lambda_{c}=0.49$, the system behaves like a topological insulator. Beyond the critical $\lambda_{c}$ (denoted by solid red), it becomes a trivial insulator with $\sigma_{xy}=0$ (that is, $C=0$). Similar observation is noted in Fig. 16$(c)$ for $\alpha=0.7$ where the plateaus at values $|C|e^{2}/h$ exist for $\lambda_{eph}$ values that are in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime.
@@
In a scenario where $\lambda_{eph}=0=M$, the Hall conductivity shows plateaus (but with small kinks on the plateaus due to the presence of the distorted FB) as long as the Fermi level lies in between the bulk gap Mondal and Basu (2023). However, in our case (with $\lambda_{eph}\neq 0$ and $M\neq 0$), the nature of the Hall conductivity deviates significantly as there exists a cumulative effect arising from the interplay of the three parameters, namely the Fermi energy ($E_{f}$), the mass term ($M$) and the e-ph coupling strength ($\lambda_{eph}$). The reason for the plateaus to become tinier can be explained with the help of the bulk spectra, which are mainly affected by $M$ and $\lambda_{eph}$ for different $\alpha$ values. It is understandable that a significant width of the plateau is dependent on how accurately we fix the Fermi level in the bulk gap. As discussed in Sec. III.1, the individual bulk bands shrink due to the Holstein factor (Eq. (8)), and the whole band structure shifts vertically down by the polaron shift energy ($\lambda_{eph}^{2}\hbar\omega_{0}$). Due to the band narrowing caused by the Holstein factor, the gap between the distorted (because of the Haldane term, $\lambda$) FB and VB decreases, makes it difficult for the Fermi level to lie ‘properly’ in between the bulk gap, making the plateaus less prominent, especially for higher values of $\lambda_{eph}$. Additionally, we find that the increase in $\sigma_{xy}$ as a function of $E_{f}$ can be explained as follows. As observed in Sec. III.1, $M$ breaks the valley degeneracy and the interplay between $M$ and $\lambda_{eph}$ renders contrasting behaviour of the bulk bands at two valleys, that is, well-gapped at one valley and almost gapless at the other for $\lambda_{eph}<\lambda_{c}$. As $E_{f}$ is increased, it is possible that at one valley, $E_{f}$ may lie well in the gap, while it may lie in the CB as well at the other valley, which will contribute to higher $\sigma_{xy}$. The unusual behaviour of the Hall conductivity due to the presence of a distorted FB has also been reported by Singh et al. Singh and Sharma (2023a) (in the absence of e-ph coupling). Certainly, all of the above discussions become unimportant for $\lambda_{eph}>\lambda_{c}$.